Convert each of the following OH- concentrations to pH.

1.9 M

The number of significant digits is set to 4; the tolerance is +/-4%

LINK TO TEXT

2.7 µM

The number of significant digits is set to 3; the tolerance is +/-4%

LINK TO TEXT

0.039 M

The number of significant digits is set to 4; the tolerance is +/-4%

LINK TO TEXT

0.00041 M

The number of significant digits is set to 4; the tolerance is +/-4%
Answer #1

The objective is to calculate the pH to the following data. Ans Given that, pH = -log Cht] and port = -log[OH-] also, pH + PO110) РОН -og COH] E-log(0.039) = 1.41 pH = 14-poh 14-1841 pH = 12.59 iv) poH = -log[or] =-log(0.00041) 53.39 pH = 14-poh -14-